A 12% salt solution is to be mixed with a 7% salt solution to obtain 15 liters of a 10% salt solution. How much of each solution should be used?

Answer by checkley79(3341)

.12X+.07(15-X)=.10*15
.12X+1.05-.07X=1.50
.05X=1.5-1.05
.05X=.45
X=.45/.05
X=9 LITERS OF 12% SOLUTION IS USED.
15-9=6 LITERS OF 7% SOLUTION IS USED.
PROOF:
.12*9+.07*6=.10*15
1.08+.42=1.5
1.5=1.5